OpenStudy (anonymous):

hi good evening.Can someone explain to me how to solve the limit with trigonometric functions?pls. help me.....

OpenStudy (anonymous):

how to evaluate this limit. Limit 1-cos 4x/1-cos 2x as x approaches 0?

OpenStudy (anonymous):

Hint:\[\frac{1-\text{Cos}[4 x]}{1-\text{Cos}[2 x]}=\frac{2 \text{Sin}\left[\frac{\pi }{4}-2 x\right] \text{Sin}\left[\frac{\pi }{4}+2 x\right]}{2 \text{Sin}\left[\frac{\pi }{4}-x\right] \text{Sin}\left[\frac{\pi }{4}+x\right]} \]